OpenStudy (anonymous):

What is the distance between the points (8, 3) and (8, –6)? A. 3 units B. 8 units C. 9 units D. 17 units I thought the answer was 10 but that answer choice isnt here, please help!

OpenStudy (anonymous):

\[d = \sqrt{(x _{2}-x _{1})^{2}+(y _{2}-y _{1})^{2}}\] plug and chug

OpenStudy (anonymous):

I did all of that but i got 10! @iambatman

OpenStudy (anonymous):

cool username btw

OpenStudy (anonymous):

\[d = \sqrt{(8-8)^2+(-6-3)^2} \implies \sqrt{(0)^2+(9)^2} \implies \sqrt{81} = ?\]

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

Sorry it should be -9 in there, but the result is the same ^.^

OpenStudy (anonymous):

Oh wait its 9 isnt it? I know what i did wrong but could you help with another one?

OpenStudy (anonymous):

Yes 9 is right!

OpenStudy (anonymous):

Well nevermind but thanks for the help!